OpenStudy (anonymous):

how do I find the use the linearization to approximate the value of f(1.4)? I know the value of a =1 and f(a) = 1 and the approximation is an underestimate

OpenStudy (anonymous):

\[ L(x) -f(a) = f'(a)(x-a) \]

OpenStudy (anonymous):

For me it is relatively easy to remember \[ f'(a) = \frac{L(x)-f(a)}{x-a} \]

OpenStudy (anonymous):

At the very least you could say \[ f'(1)=\frac{L(x)-1}{0.4} \]

OpenStudy (anonymous):

Whoops, should have said: \[ f'(1) = \frac{L(1.4)-1}{0.4} \]

OpenStudy (anonymous):